# DeFi Arbitrage in Hedged Liquidity Tokens

### v3

@@ Abstract
Empirically, the prevailing market prices for liquidity tokens of the constant product market maker (CPMM) – as offered in practice by companies such as Uniswap – readily permit arbitrage opportunities by delta hedging the risk of the position. Herein, we investigate this arbitrage opportunity by treating the liquidity token as a derivative position in the prices of the underlying assets for the CPMM. In doing so, not dissimilar to the Black-Scholes result, we deduce risk-neutral pricing and hedging formulas for these liquidity tokens. Furthermore, with our novel pricing formula, we construct a method to calibrate a volatility to data which provides an updated (non-market) price which would not permit arbitrage if quoted by the CPMM. We conclude with a discussion of novel AMM designs which would bring the pricing of liquidity tokens into the modern financial era.
@@ 1 Introduction
We take the view that the investment of a LP in an AMM (i.e., purchasing a liquidity token) is a path-dependent perpetual derivative which earns a dividend stream (i.e., fees) based on the executed swaps on the AMM. Immediately, with this viewpoint of the liquidity token, the prevailing pricing structure quoted in the CPMM smart contract reveals itself to be based in the pre-Black-Scholes world. For example, when applied to Uniswap data, arbitrage opportunities are readily available to a sophisticated investor (see Example 2.3 below). [^1] In contrast to, e.g., [12] we consider the accounting profits/losses of the liquidity position rather than in relation to an opportunity cost (the loss-versus-rebalancing in the cited paper). Interestingly, even with the potential for arbitrage profits, when compared to the buy-and-hold strategy of the initial liquidity position, nearly 50% of LPs lose money on Uniswap [19]. As LPs form the backbone of these decentralized exchanges, these losses emphasize the need to introduce hedging strategies for liquidity tokens; if investors were to withdraw liquidity en masse (due to the high risk of the investment) then the entire DeFi paradigm would fail its primary task to act as a financial intermediary. The goal of this work is to construct risk-neutral pricing and hedging theory for liquidity tokens in the CPMM to bring this DeFi product into the modern financial world.
@@
Our primary contributions and innovations for the pricing and hedging of the liquidity position in a CPMM are threefold. First, in treating this liquidity position as a derivative on the underlying assets, we find the optimal execution of the position. That is, we deduce conditions for when a risk-neutral investor would optimally invest in (or withdraw from) the CPMM as a LP. With this optimal execution, second, we are able to produce a risk-neutral valuation for a liquidity token. As a direct consequence, the Greeks and hedging strategies for this position can be readily constructed. As far as the authors are aware, a formal discussion of the Greeks of the liquidity token has never been undertaken previously. Notably, as nearly 50% of LPs lose money on Uniswap [19], the introduction of a hedging strategy is of vital importance. Third, we bring our pricing and hedging theory to data in order to understand its performance in practice. We find that the prevailing market price for the CPMM liquidity token readily admits arbitrage opportunities that investors can exploit. Bringing the theory to the data, we construct a calibrated arbitrage-free price for the liquidity token.
@@ 2 Motivating Example
The primary motivation of this paper is to understand how to hedge an investment into the constant product market maker (CPMM). Within Section 2.1, the basic construction of an AMM – and specifically a CPMM – is provided. With these details, in Section 2.2, we consider the value of the CPMM and its delta hedging position when being priced at the current market rate. This valuation is provided using Uniswap data to demonstrate that arbitrage opportunities exist in the current market setup. The subsequent sections of this work focus on updating the formulas for pricing and hedging so as to properly eliminate arbitrage opportunities.
@@ 2.1 Background on Automated Market Maker
For the remainder of this work, we will consider the constant product market maker (CPMM) in the $n=2$ asset setting as is utilized by Uniswap and Sushiswap pools (i.e., $u(x,y):=xy$). In addition, throughout, we take the second asset as the numéraire (i.e., $P(x,y):=P_{1}^{2}(x,y)$).
@@ 2.2 Hedging a CPMM
Given this value of the liquidity token, we can eliminate entirely the riskiness of the position by trading the underlying token appropriately (see, e.g., the discussion in and preceding [16, Chapter 2.4]). Specifically, we can hedge this position by holding $\Delta$ units of the underlying where $\Delta$ is the sensitivity of the position to the token price, i.e., $\Delta=\frac{d}{dP}2\sqrt{P}=1/\sqrt{P}$. In theory, by purchasing and rebalancing this position, we expect to perfectly hedge the position; however, this hedge may fail during periods of extreme volatility as delta hedging only provides protections along the linearly approximated portfolio. In the following numerical example, we take data on a Uniswap pool in order to look for any arbitrage opportunities, i.e., for trends in the hedged position. Surprisingly, given the theory on delta hedging, we find opportunities for risk-free profits with frequent hedges. To preview future results, in the next section we provide a theory of risk-neutral pricing of a liquidity token and deduce the Greeks of this liquidity position; these results are summarized in Table 1. Herein, we consider the hedging and rebalancing strategy under negligible fees as is taken in the Black-Scholes framework.
@@
Consider a USDC/WETH Uniswap v3 pool on the Polygon blockchain ($\Delta t=2$ seconds) with $\gamma=5bps$ fees between market creation (December 20, 2021) and June 21, 2025 with the investment made over the entire price line to mimic the CPMM. [^2] This data is taken from smart contract 0x45dda9cb7c25131df268515131f647d726f50608. This pool was chosen as it is a representative Uniswap v3 pool with high liquidity. For this example, we will assume that the risk-free rate $r=0\%$ throughout the period of study for simplicity and to match the rate earned on USDC. In Figure 1(a), we show the discounted values of both a 1 USDC liquidity position (i.e., $\sqrt{P}/\sqrt{P_{0}}$ plus the collected fees) and of the delta hedged positions (with rebalancing of the delta hedge at different frequencies ranging from every block to every day). As expected, the volatility of the hedged position is significantly lower than that of the unhedged position. However, as made clear in Figure 1(b), when delta is hedged every block, then the hedged position has a distinct positive trend line indicating the presence of an arbitrage opportunity except on distinct dates in which the hedging fails due to large intra-block price swings. We hypothesize that a better hedging strategy (e.g., Gamma hedging) is needed to reduce these jumps. Notably, from February 2022 to March 2024, there is a nearly unbroken stretch of profitable arbitrage opportunities by implementing this delta hedging strategy. As we reduce the hedging frequency, we find a positive trend exists for rehedging every minute, though the profitability disappears for hedging every hour or day.
@@ 3.1 Market Model
Following Assumption 3.1, the measure $\mathbb{P}$ is a risk-neutral measure. In fact, due to the completeness of the constructed market, $\mathbb{P}$ is the unique risk-neutral measure. [^3] We refer the interested reader to [14, 10] for the discussion on the Fundamental Theorems of Asset Pricing and to [16, Chapters 1 and 2] that show the existence of a hedge in a complete market driven by a Brownian Motion as well as pricing by replication in such a market. Throughout the remainder of this work, we will exploit this fact in order to consider the pricing and hedging of a liquidity token in a CPMM.
@@ 3.3 Greeks
As we can describe the value of the liquidity position in the constant function market maker via Theorem 3.2 and Corollary 3.3, it is valuable also to understand how to hedge the risks of this position. For this purpose we will consider various Greeks for the liquidity token. Herein we will focus specifically on the Greeks at block times though, utilizing the forms of Corollary 3.3, the Greeks can also be computed between blocks.
@@ 4.1 Implied Volatility
Consider a constant product market maker on the Polygon blockchain ($\Delta t=2$ seconds). For these examples, recall that the risk-neutral valuation and other formulas employed throughout this work utilize the realized fee level $\hat{\gamma}=\gamma/(1-\gamma)$. For any choice of $\gamma$, there are two possible outcomes. The less interesting possibility is that $V_{0}(P)<2\sqrt{P}$ so that depositing liquidity for the price of $2\sqrt{P}$ is expected to lose value, i.e., the expected discounted value of the fees would not cover this initial deposit value. [^4] Within this work, we have assumed that it is not possible to sell liquidity tokens short and, therefore, no arbitrage opportunity exists if $V_{0}(P)<2\sqrt{P}$. The more interesting scenario is where $V_{0}(P)\geq 2\sqrt{P}$ which (with a strict inequality) results in an arbitrage opportunity since the purchase price of the liquidity tokens $2\sqrt{P}$ is below the risk-neutral valuation of the fees. Therefore, in this latter scenario, a rational investor can deposit the liquidity for $2\sqrt{P}$ and, with appropriate hedging, obtain the value $V_{0}(P)$. [^5] We refer the interested reader to [16, Chapter 1.4] for an initial discussion on arbitrage.
@@ 5 Discussion
Given that we found that the prevailing market prices for CPMMs exhibit arbitrage opportunities, it is important to construct new AMM designs that permit freely floating pricing for liquidity tokens. That is, where the price of a liquidity token depends explicitly on the number of outstanding tokens which does not exist at present. This would also permit, e.g., varying implied volatilities instead of the flat implied volatility based on the current pricing scheme (as given in Section 4.1).
@@
Following the mispricing identified in Example 4.5, introducing variable minting or burning values of the next (marginal) liquidity token $V(P,L)=2v(L)\sqrt{P}$ for some strictly increasing function $v:\mathbb{R}_{++}\to\mathbb{R}_{++}$ of the outstanding liquidity tokens $L>0$. To recreate such a structure, the CPMM structure needs to be updated. Specifically, taking $xy=\ell(L)$ – generalizing $xy=L^{2}$ as taken for CPMMs at present – for $\ell:\mathbb{R}_{++}\to\mathbb{R}_{++}$ twice continuously differentiable with $\ell^{\prime}(L)>0$ and $2\ell(L)\ell^{\prime\prime}(L)>\ell^{\prime}(L)^{2}$ for every $L>0$ results in the updated pricing $v(L)=\frac{\ell^{\prime}(L)}{2\sqrt{\ell(L)}}$. Note that this updated pricing, generally, depends explicitly on the outstanding liquidity tokens $L$. For example, inspired by the classical quadratic approach to liquidity tokens, taking $\ell(L)=L^{2\alpha}$ for $\alpha>1$ results in the updated pricing scheme with $v(L)=\alpha L^{\alpha-1}$. [^7] A mathematical treatment of varying the price of liquidity tokens following $\ell(L)=L^{2\alpha}$ is presented in [5]. With this updated pricing, the price of liquidity tokens will fluctuate as the total amount of market liquidity changes. Akin to traditional derivatives markets, LPs can use the implied volatility (so that $V_{0}(P;\sigma)=2v(L)\sqrt{P}$ to generalize the discussion of Section 4.1) in order to evaluate the performance of their investment. However, since other CPMMs exist in the market, a savvy investor would never pay more than $2\sqrt{P}$ for this contract. Therefore, though in theory having fully variable pricing is possible, it would never succeed unless, and until, all CPMMs allow for variable pricing.
@@ 6 Conclusion
Evaluating data from a CPMM indicates that arbitrage opportunities readily exist based on the prevailing market pricing for a liquidity token from, e.g., Uniswap. The constant pricing scheme considered in practice can be viewed akin to the pre-Black-Scholes world for derivatives. Within this work we have determined a risk-neutral pricing theory for CPMMs. With this theory we have revisited the data to determine an approximating arbitrage-free price. Furthermore, we propose two novel AMM designs so that the pricing of the liquidity tokens are variable in time based on the demand for such tokens.
Though we focused solely on the CPMM within this work, we conjecture similar arbitrage opportunities can be found in other AMM designs. In particular, we wish to highlight the concentrated liquidity designs of Uniswap v3; we conjecture that the optimal investment strategy for the concentrated liquidity would include finite stopping times which may complicate the risk-neutral pricing. In comparison to the flat implied volatility curve for the CPMM (see Section 4.1), the concentrated liquidity structure would permit a DeFi implied volatility curve; such a structure could be of great interest for practitioners to more accurately price and hedge DeFi risks. We believe a study of such constructions would be of great interest.

### v4

# A Derivative Pricing Perspective on Liquidity Tokens in Constant Product Market Makers

@@ Abstract
In decentralized finance, any individual can pool their assets into an automated market maker (AMM) – herein we focus on the constant product market maker (CPMM) – in exchange for a claim on a fraction of future pool assets and fees earned from the market making operations. This position is represented by a liquidity token, whose prevailing on-chain price is effectively the initial deposited assets. Though this price is well-defined, we treat the liquidity token as a derivative position in the prices of the underlying assets for the CPMM in order to deduce risk-neutral pricing and hedging formulas, not dissimilar to the Black-Scholes result. Adopting this perspective, in a frictionless environment, hedging the CPMM liquidity token under fair valuation should produce a riskless process, which therefore grows at the risk-free rate, something that is not seen in empirical case studies under the prevailing price. With our novel pricing formula, we construct a method to calibrate a volatility to data which provides an updated (non-market) valuation which is consistent with the (near-continuous) replication strategy out-of-sample. We conclude with a discussion of novel AMM design considerations motivated by this derivative-pricing perspective.
@@ 1 Introduction
We take the view that the investment of a LP in an AMM (i.e., purchasing a liquidity token) is a path-dependent perpetual derivative which earns a dividend stream (i.e., fees) based on the executed swaps on the AMM and its market price is the cost of the liquidity deposited. Immediately, with this viewpoint of the liquidity token, the prevailing pricing structure quoted in the CPMM smart contract reveals itself to be based in the pre-Black-Scholes world. For example, when applied to Uniswap data, arbitrage opportunities arise under self-financing hedging strategies (see Example 2.3 below). [^1] Throughout this work, we consider arbitrage only in a frictionless environment. [^2] In contrast to, e.g., [11] we consider the accounting profits/losses of the liquidity position rather than in relation to an opportunity cost (the loss-versus-rebalancing in the cited paper). Interestingly, even with the potential for arbitrage profits, when compared to the buy-and-hold strategy of the initial liquidity position, nearly 50% of LPs lose money on Uniswap [18]. As LPs form the backbone of these decentralized exchanges, these losses emphasize the need to introduce hedging strategies for liquidity tokens; if investors were to withdraw liquidity en masse (due to the high risk of the investment) then the entire DeFi paradigm would fail its primary task to act as a financial intermediary. The goal of this work is to construct risk-neutral pricing and hedging theory for liquidity tokens in the CPMM to bring this DeFi product into the modern financial world.
@@
Our primary contributions and innovations for the pricing and hedging of the liquidity position in a CPMM are threefold. First, in treating this liquidity position as a derivative on the underlying assets, we find the optimal execution of the position. That is, we deduce conditions for when a risk-neutral investor would optimally invest in (or withdraw from) the CPMM as a LP. With this optimal execution, second, we are able to produce a risk-neutral valuation for a liquidity token. As a direct consequence, the Greeks and hedging strategies for this position can be readily constructed. As far as the authors are aware, a formal discussion of the Greeks of the liquidity token has never been undertaken previously. Notably, as nearly 50% of LPs lose money on Uniswap [18], the introduction of a hedging strategy is of vital importance. Third, we bring our pricing and hedging theory to data in order to understand its performance in practice. We find that the prevailing market price for the CPMM liquidity token is inconsistent with the cost of the replicating portfolio in a frictionless environment with frequent rebalancing. Bringing the theory to the data, we construct a calibrated arbitrage-free price for the liquidity token.
@@ 2 Motivating Example
The primary motivation of this paper is to understand how to hedge an investment into the constant product market maker (CPMM). Within Section 2.1, the basic construction of an automated market maker (AMM) – and specifically a CPMM – is provided. With these details, in Section 2.2, we consider the value of the CPMM and its delta hedging position when being priced at the current market rate. This valuation is provided using Uniswap data to demonstrate that arbitrage opportunities exist in the current market setup. The subsequent sections of this work focus on updating the formulas for pricing and hedging so as to properly eliminate arbitrage opportunities.
@@ 2.1 Background on Automated Market Maker
For the remainder of this work, we will consider the CPMM in the $n=2$ asset setting as is utilized by Uniswap and Sushiswap pools (i.e., $u(x,y):=xy$). In addition, throughout, we take the second asset as the numéraire (i.e., $P(x,y):=P_{1}^{2}(x,y)$).
@@ 2.2 Hedging a CPMM
Given this value of the liquidity token, we can eliminate entirely the riskiness of the position by trading the underlying token appropriately (see, e.g., the discussion in and preceding [15, Chapter 2.4]). Specifically, we can hedge this position by holding $\Delta$ units of the underlying where $\Delta$ is the sensitivity of the position to the token price, i.e., $\Delta=\frac{d}{dP}2\sqrt{P}=1/\sqrt{P}$. In theory, by purchasing and rebalancing this position, we expect to perfectly hedge the liquidity token; however, this hedge may fail during periods of extreme volatility as delta hedging only provides protections along the linearly approximated portfolio. In the following numerical example, we take data on a Uniswap pool in order to look for any arbitrage opportunities, i.e., for trends in the hedged position. Surprisingly, given the theory on delta hedging, we observe a persistent positive drift in the discounted, frequently-hedged position. To preview future results, in the next section we provide a theory of risk-neutral pricing of a liquidity token and deduce the Greeks of this liquidity position; these results are summarized in Table 1. Herein, we consider the hedging and rebalancing strategy under negligible fees as is taken in the Black-Scholes framework.
@@
Consider a USDC/WETH Uniswap v3 pool on the Polygon blockchain ($\Delta t=2$ seconds) with $\gamma=5bps$ fees between market creation (December 20, 2021) and June 21, 2025 with the investment made over the entire price line to mimic the CPMM. [^3] This data is taken from smart contract 0x45dda9cb7c25131df268515131f647d726f50608. This pool was chosen as it is a representative Uniswap v3 pool with high liquidity. For this example, we will assume that the risk-free rate $r=0\%$ throughout the period of study for simplicity and to match the rate earned on USDC. In Figure 1(a), we show the discounted values of both a 1 USDC liquidity position (i.e., $\sqrt{P}/\sqrt{P_{0}}$ plus the collected fees) and of the delta hedged positions (with rebalancing of the delta hedge at different frequencies ranging from every block to every day). As expected, the volatility of the hedged position is significantly lower than that of the unhedged position. However, as made clear in Figure 1(b), when delta is hedged every block, then the hedged position has a distinct positive trend line indicating the presence of an arbitrage opportunity except on distinct dates in which the hedging fails due to large intra-block price swings. We hypothesize that a better hedging strategy (e.g., Gamma hedging) is needed to reduce these jumps. Notably, from February 2022 to March 2024, there is a nearly unbroken stretch of profitable arbitrage opportunities by implementing this delta hedging strategy. As we reduce the hedging frequency, we find a positive trend exists for rehedging every minute as well.
@@ 3.1 Market Model
Following Assumption 3.1, the measure $\mathbb{P}$ is a risk-neutral measure. In fact, due to the completeness of the constructed market, $\mathbb{P}$ is the unique risk-neutral measure. [^4] We refer the interested reader to [13, 9] for the discussion on the Fundamental Theorems of Asset Pricing and to [15, Chapters 1 and 2] that show the existence of a hedge in a complete market driven by a Brownian Motion as well as pricing by replication in such a market. Throughout the remainder of this work, we will exploit this fact in order to consider the pricing and hedging of a liquidity token in a CPMM.
@@ 3.3 Greeks
As we can describe the value of the liquidity position in the CPMM via Theorem 3.2 and Corollary 3.3, it is valuable also to understand how to hedge the risks of this position. For this purpose we will consider various Greeks for the liquidity token. Herein we will focus specifically on the Greeks at block times though, utilizing the forms of Corollary 3.3, the Greeks can also be computed between blocks.
@@ 4.1 Implied Volatility
Consider a CPMM on the Polygon blockchain ($\Delta t=2$ seconds). For these examples, recall that the risk-neutral valuation and other formulas employed throughout this work utilize the realized fee level $\hat{\gamma}=\gamma/(1-\gamma)$. For any choice of $\gamma$, there are two possible outcomes. The less interesting possibility is that $V_{0}(P)<2\sqrt{P}$ so that depositing liquidity for the price of $2\sqrt{P}$ is expected to lose value, i.e., the expected discounted value of the fees would not cover this initial deposit value. [^5] Within this work, we have assumed that it is not possible to sell liquidity tokens short and, therefore, no arbitrage opportunity exists if $V_{0}(P)<2\sqrt{P}$. The more interesting scenario is where $V_{0}(P)\geq 2\sqrt{P}$ which (with a strict inequality) results in an arbitrage opportunity since the purchase price of the liquidity tokens $2\sqrt{P}$ is below the risk-neutral valuation of the fees. Therefore, in this latter scenario, a rational investor can deposit the liquidity for $2\sqrt{P}$ and, with appropriate hedging, obtain the value $V_{0}(P)$. [^6] We refer the interested reader to [15, Chapter 1.4] for an initial discussion on arbitrage.
@@ 5 Discussion
Given that we found that the prevailing market prices for CPMMs can exhibit arbitrage opportunities, it is important to construct new AMM designs that permit freely floating pricing for liquidity tokens. That is, where the price of a liquidity token depends explicitly on the number of outstanding tokens which does not exist at present. This would also permit, e.g., varying implied volatilities instead of the flat implied volatility based on the current pricing scheme (as given in Section 4.1).
@@
Following the mispricing identified in Example 4.5, introducing variable minting or burning values of the next (marginal) liquidity token $V(P,L)=2v(L)\sqrt{P}$ for some strictly increasing function $v:\mathbb{R}_{++}\to\mathbb{R}_{++}$ of the outstanding liquidity tokens $L>0$. To recreate such a structure, the CPMM structure needs to be updated. Specifically, taking $xy=\ell(L)$ – generalizing $xy=L^{2}$ as taken for CPMMs at present – for $\ell:\mathbb{R}_{++}\to\mathbb{R}_{++}$ twice continuously differentiable with $\ell^{\prime}(L)>0$ and $2\ell(L)\ell^{\prime\prime}(L)>\ell^{\prime}(L)^{2}$ for every $L>0$ results in the updated pricing $v(L)=\frac{\ell^{\prime}(L)}{2\sqrt{\ell(L)}}$. Note that this updated pricing, generally, depends explicitly on the outstanding liquidity tokens $L$. For example, inspired by the classical quadratic approach to liquidity tokens, taking $\ell(L)=L^{2\alpha}$ for $\alpha>1$ results in the updated pricing scheme with $v(L)=\alpha L^{\alpha-1}$. With this updated pricing, the price of liquidity tokens will fluctuate as the total amount of market liquidity changes. Akin to traditional derivatives markets, LPs can use the implied volatility (so that $V_{0}(P;\sigma)=2v(L)\sqrt{P}$ to generalize the discussion of Section 4.1) in order to evaluate the performance of their investment. However, since other CPMMs exist in the market, a savvy investor would never pay more than $2\sqrt{P}$ for this contract. Therefore, though in theory having fully variable pricing is possible, it would never succeed unless, and until, all CPMMs allow for variable pricing.
@@ 6 Conclusion
Evaluating data from a CPMM indicates that hedging the liquidity token from, e.g., Uniswap, can result in arbitrage opportunities. The constant pricing scheme considered in practice can be viewed akin to the pre-Black-Scholes world for derivatives. Within this work we have determined a risk-neutral pricing theory for CPMMs. With this theory we have revisited the data to determine an approximating arbitrage-free price. Furthermore, we propose two novel AMM designs so that the pricing of the liquidity tokens are variable in time based on the demand for such tokens.
Though we focused solely on the CPMM within this work, we conjecture similar mispricing can be found in other AMM designs. In particular, we wish to highlight the concentrated liquidity designs of Uniswap v3; we conjecture that the optimal investment strategy for the concentrated liquidity would include finite stopping times which may complicate the risk-neutral pricing. In comparison to the flat implied volatility curve for the CPMM (see Section 4.1), the concentrated liquidity structure would permit a DeFi implied volatility curve; such a structure could be of great interest for practitioners to more accurately price and hedge DeFi risks. We believe a study of such constructions would be of great interest.